# A convergent adaptive finite element stochastic Galerkin method based on multilevel expansions of random fields

### v1

@@ 3.2. Refining the triangulations
that outputs the componentwise smallest sequence of meshes such that for each $\nu\in F$, $\tilde{\mathcal{T}}_{\nu}$ is conforming and $\operatorname{span}\{\psi_{\lambda}\}_{\lambda\in\Theta_{\nu}}\subseteq V(\tilde{\mathcal{T}}_{\nu})$. Note that in a setting where $\Lambda^{0}\subseteq\Lambda$, the meshes $\tilde{\mathcal{T}}_{\nu}$ are refinements of the initial meshes $\mathcal{T}_{\nu}$ that are given by the frame elements in $\Lambda^{0}$: $(\mathcal{T}_{\nu})_{\nu\in F^{0}}=\text{{Mesh}}(\Lambda^{0})$.
@@ 3.3. Solving Galerkin systems
The condition (3.16) can be verified directly in the PCG method as follows. Let $\mathbf{A}\in\mathbb{R}^{N\times N}$ with $N=\dim\mathcal{V}({\mathbb{T}})$ be the coefficient matrix such that for nodal coordinate vectors $\mathbf{u},\mathbf{v}\in\mathbb{R}^{N}$ of functions $u,v\in\mathcal{V}({\mathbb{T}})$ we have $\langle\mathbf{A}\mathbf{u},\mathbf{v}\rangle=\int_{Y}\int_{D}\nabla u(x,y)\cdot\nabla v(x,y)\,\mathrm{d}x\,\mathrm{d}y$. We use the modification of the BPX preconditioner to general meshes generated by newest vertex bisection analyzed in [13]. For this preconditioner with matrix representation $\mathbf{P}$, there exist $c_{p},C_{P}>0$ such that we have the spectral equivalence
@@
The number of steps of the PCG method in Algorithm 5 terminates after a uniform finite number of steps since we require a fixed relative accuracy. Hence, the computational complexity of the algorithm does not deteriorate when successive mesh refinements take place.
@@ 4. Convergence analysis
(4.3) $0<(C+2)\zeta<1\qquad\text{with}\qquad C^{2}=\frac{C_{\Psi}^{2}C_{B}C_{B,\Psi}}{c_{\Psi}^{2}c_{B}}$
@@
(4.4) $\omega_{0}>(C+1)\frac{\zeta}{1-\zeta}.$
@@
are equivalent and thus error reduction is achieved due to (4.2) in Lemma 4.2. The terms $2^{-\alpha\ell}$, $\zeta_{\hat{J}}$ and $\rho$ appearing in $\gamma(\zeta,\ell,\rho,\hat{J})$ from 3.12 can be made arbitrarily small for sufficiently large $\hat{J}$ and $\ell$ and sufficiently small $\rho$. Thus the condition (4.3) is satisfiable whenever
@@ 5. Numerical Experiments
Figure 3. Computed residual bounds for $d=1$ as a function of total number of degrees of freedom $\dim\mathcal{V}({\mathbb{T}})\eqsim N(\mathbb{T})$ of the current approximation of $u$ (solid grey lines), degrees of stochastic freedom #$\#F$ (solid black lines), and elapsed computation time in seconds (dash-dotted line).
Figure 4. Computed residual bounds for $d=2$ as a function of total number of degrees of freedom $\dim\mathcal{V}({\mathbb{T}})\eqsim N(\mathbb{T})$ of the current approximation of $u$ (solid grey lines), number of product Legendre polynomials #$\#F$ (solid black lines), and elapsed computation time in seconds (dash-dotted line).

### v2

@@ 3.2. Refining the triangulations
that outputs the componentwise smallest sequence of meshes such that for each $\nu\in F$, $\tilde{\mathcal{T}}_{\nu}$ is conforming and $\operatorname{span}\{\psi_{\lambda}\}_{\lambda\in\Theta_{\nu}}\subseteq V(\tilde{\mathcal{T}}_{\nu})$. Note that in a setting where $\Lambda^{0}\subseteq\Lambda$, the meshes $\tilde{\mathcal{T}}_{\nu}$ are refinements of the initial meshes $\mathcal{T}_{\nu}$ that are given by the frame elements in $\Lambda^{0}$, that is, $(\mathcal{T}_{\nu})_{\nu\in F^{0}}=\text{{Mesh}}(\Lambda^{0})$.
@@ 3.3. Solving Galerkin systems
The condition (3.17) can be verified directly in the PCG method as follows. Let $\mathbf{A}\in\mathbb{R}^{N\times N}$ with $N=\dim\mathcal{V}({\mathbb{T}})$ be the coefficient matrix such that for nodal coordinate vectors $\mathbf{u},\mathbf{v}\in\mathbb{R}^{N}$ of functions $u,v\in\mathcal{V}({\mathbb{T}})$ we have $\langle\mathbf{A}\mathbf{u},\mathbf{v}\rangle=\int_{Y}\int_{D}\nabla u(x,y)\cdot\nabla v(x,y)\,\mathrm{d}x\,\mathrm{d}y$. We use the modification of the BPX preconditioner to general meshes generated by newest vertex bisection analyzed in [13]; an alternative multigrid approach is presented in [29]. For this preconditioner with matrix representation $\mathbf{P}$, there exist $c_{p},C_{P}>0$ such that we have the spectral equivalence
@@
The number of steps of the PCG method in Algorithm 5 terminates after a uniform finite number of steps since we require a fixed relative accuracy. Hence, the computational complexity of the algorithm does not deteriorate when successive mesh refinements take place. Furthermore, the operator compression parameter $\ell$ is chosen as a fixed value, similarly depending only on a fixed relative tolerance. The complexity of one iteration of the CG method can thus be bounded by a fixed multiple of $N({\mathbb{T}}^{k})$, and consequently the same holds true for each call of GalerkinSolve in Algorithm 6.
@@ 4. Convergence analysis
(4.3) $0<(C_{B,\Psi}+2)\zeta<1\qquad\text{with}\qquad C_{B,\Psi}=\frac{C_{\Psi}^{2}C_{B}}{c_{\Psi}^{2}c_{B}}$
@@
(4.4) $\omega_{0}>(C_{B,\Psi}+1)\frac{\zeta}{1-\zeta}.$
@@
are equivalent and thus error reduction is achieved due to (4.2) in Lemma 4.2. The terms $2^{-\alpha\ell}$, $\zeta_{\hat{J}}$ and $\rho$ appearing in $\gamma(\zeta,\ell,\rho,\hat{J})$ from 3.12 can be made arbitrarily small for sufficiently large $\hat{J}$ and $\ell$ and sufficiently small $\rho$. Thus the condition (4.3) is achievable whenever
@@
Remark 4.5.
The convergence estimate of Theorem 4.3 can be shown in the same manner without requiring the functions $\theta_{\mu}$ in the expansion of the diffusion cofficient to have wavelet-type multilevel localization properties. For enforcing the uniform error reduction based on our construction, only piecewise polynomial structure (or approximation) on some uniformly refined mesh of each function $\theta_{\mu}$ in the expansion of the diffusion coefficient is required, which also covers KL-type expansions. Specifically, as noted in Remark 2.3(c), 2 can be replaced by only requiring that for a fixed $m$, for each $\mu\in{\mathcal{M}}$ there exists $j(\mu)$ such that $\theta_{\mu}\in\mathbb{P}_{m}(\hat{\mathcal{T}}_{j(\mu)})$. The semidiscrete operator compression in 3.1 utilizes the multilevel structure of $(\theta_{\mu})_{\mu\in{\mathcal{M}}}$, but a similar compression for KL expansions using summability of the norms $\lVert\theta_{\mu}\rVert_{L_{\infty}}$ was shown in [21]; note, however, that this generally leads to suboptimal complexity. The required level of additional refinements in Lemma 2.4 is still bounded uniformly in terms of $m$, independently of $j(\mu)$. The error bound of 3.7 can again be shown in the same manner using $j(\mu)$. Thus, the convergence result Theorem 4.3 follows. However, our method is designed to exploit the stronger 2 to obtain the computational complexity bounds summarized in Remark 4.4.
@@ 5. Numerical Experiments
Figure 3. Computed residual bounds for $d=1$ as a function of total number of degrees of freedom $\dim\mathcal{V}({\mathbb{T}})\eqsim N(\mathbb{T})$ of the current approximation of $u$ (solid gray lines), degrees of stochastic freedom #$\#F$ (solid black lines), and elapsed computation time in seconds (dash-dotted line).
Figure 4. Computed residual bounds for $d=2$ as a function of total number of degrees of freedom $\dim\mathcal{V}({\mathbb{T}})\eqsim N(\mathbb{T})$ of the current approximation of $u$ (solid gray lines), number of product Legendre polynomials #$\#F$ (solid black lines), and elapsed computation time in seconds (dash-dotted line).